# A Mapper algorithm with implicit intervals and its optimization

### v1

@@ Abstract
The Mapper algorithm is an essential tool for visualizing complex, high dimensional data in topology data analysis (TDA) and has been widely used in biomedical research. It outputs a combinatorial graph whose structure implies the shape of the data. However,the need for manual parameter tuning and fixed intervals, along with fixed overlapping ratios may impede the performance of the standard Mapper algorithm. Variants of the standard Mapper algorithms have been developed to address these limitations, yet most of them still require manual tuning of parameters. Additionally, many of these variants, including the standard version found in the literature, were built within a deterministic framework and overlooked the uncertainty inherent in the data. To relax these limitations, in this work, we introduce a novel framework that implicitly represents intervals through a hidden assignment matrix, enabling automatic parameter optimization via stochastic gradient descent. In this work, we develop a soft Mapper framework based on a Gaussian mixture model(GMM) for flexible and implicit interval construction. We further illustrate the robustness of the soft Mapper algorithm by introducing the Mapper graph mode as a point estimation for the output graph. Moreover, a stochastic gradient descent algorithm with a specific topological loss function is proposed for optimizing parameters in the model. Both simulation and application studies demonstrate its effectiveness in capturing the underlying topological structures. In addition, the application to an RNA expression dataset obtained from the Mount Sinai/JJ Peters VA Medical Center Brain Bank (MSBB) successfully identifies a distinct subgroup of Alzheimer’s Disease. The implementation of our method is available at https://github.com/FarmerTao/Implicit-interval-Mapper.git.
@@ 1 Introduction
The Mapper algorithm is a powerful tool in topology data analysis (TDA) used to explore the shape of datasets. Initially introduced for 3D object recognition (Singh et al., 2007), the Mapper algorithm has demonstrated remarkable efficiency in extracting topological information across various data types. Its application spans various areas, notably in biomedical research (Skaf and Laubenbacher, 2022). For example, it successfully identified new breast cancer subgroups with superior survival rates (Nicolau et al., 2011). In single cell RNA-Seq analysis, a Mapper-based algorithm was used to study unbiased temporal transcriptional regulation (Rizvi et al., 2017). The variants of Mapper algorithms were also used to uncover higher-order structures of complex phenomics data (Kamruzzaman et al., 2021). Moreover, many efforts have been made to integrate the Mapper algorithm with other machine learning techniques. For instance, Bodnar et al. (2021) enhanced the performance of a graph neural network by incorporating the Mapper algorithm as a pooling operation. When combined with autoencoders, the Mapper algorithm can be used as a robust classifier (Cyranka et al., 2019), remedying the shortcomings of traditional convolutional neural networks, such as susceptibility to gradient-based attacks. Recently, a visualization method called ShapeVis which was inspired by the Mapper algorithm was proposed to produce more concise topological structure than standard Mapper graph (Kumari et al., 2020). Besides advances of the Mapper algorithms in applications, there has been some new progress in theory. Carrière et al. (2018) analyzed the Mapper algorithm from statistical perspective and introduced a novel method for parameter selection. Other studies have explored the algorithm’s convergence to Reeb graphs (Brown et al., 2021), and examined its topological properties (Dey et al., 2017).
@@
The standard Mapper algorithm generates a graph that captures the shape of the data by dividing the support of the filtered data into several overlapped intervals with a fixed length (Singh et al., 2007). The major limitations of the standard Mapper algorithm are the requirements of fixed length of intervals and overlapping rate. Many algorithms were proposed to address these limitations. For instance, F-Mapper (Bui et al., 2020) utilized the fuzzy clustering algorithm to allow flexible interval partitioning by introducing an additional threshold parameter to control the degree of overlap. However, the introduction of the parameter also increased the algorithm’s computational complexity. In a different vein, the ensemble version of Mapper was proposed by Kang and Lim (2021); Fitzpatrick et al. (2023), which can recover the topology of datasets without parameter tuning. However, this method has a significant high computational cost due to the execution of multiple instances of the Mapper algorithm. Differently, the Ball Mapper (Dłotko, 2019) directly constructed cover on a given dataset, avoiding the difficulty of choosing filter functions. Nevertheless, its construction process is somewhat arbitrary and could inadvertently introduce extra parameters, increasing computational complexity. The D-Mapper, as introduced by Tao and Ge (2024), developed a probabilistic framework to construct intervals of the projected data based on a mixture distribution. This approach offers flexibility in interval construction. Additionally, the Soft Mapper, proposed by Oulhaj et al. (2024), a smoother version of the standard Mapper algorithm. This algorithm focused on the optimization of filter functions by minimizing the topological loss.
@@ 2.1 Mapper algorithm
The Mapper algorithm has been widely used to output a graph of a given dataset to represent its topological structure (Dey et al., 2017), the output graph can be regarded as a discrete version of the Reeb graph. The core concept of the Mapper algorithm is as follows. First, it projects a given high-dimensional dataset into a lower-dimensional space. Then, it constructs a reasonable cover for the projected data and pulls back the original data points according to each interval and clustering. Finally, each cluster is treated as a node of the output graph, and an edge is added between two clusters if they share any data points.
@@
Consider a dataset with $n$ data points, $\{\boldsymbol{x_{1}},\ldots,\boldsymbol{x_{n}}\}$, $\boldsymbol{x_{i}}\in\mathbb{R}^{d}$, $i=1,\ldots,n$. Let $\mathbb{X}_{n}$ be the support of this dataset. Denote $f(\cdot):\mathbb{X}^{n}\rightarrow\mathbb{R}^{m}$, $m<n$, a single-valued filter function, which maps each data point from $\mathbb{X}^{n}$ to a lower-dimensional space, $\mathbb{R}^{m}$ ($m\geq 1$). As in most of literature Bui et al. (2020); Kang and Lim (2021); Fitzpatrick et al. (2023), here $m$ is set to $1$. We use a simple example to illustrate the Mapper algorithm. The example dataset has a cross shape, as shown in Figure 1 (a). The Mapper algorithm first projects data onto a real line using a filter function, $f(\boldsymbol{x})=y,~{}\boldsymbol{x}\in\mathbb{R}^{2},~{}y\in\mathbb{R}$, as shown in the upper pannel of Figure 1 (b), the filter function is set to $f(\boldsymbol{x})=\frac{1}{n}\sum_{i=1}^{n}||\boldsymbol{x}-\boldsymbol{x}_{i}||$, i.e. the mean distance of $\boldsymbol{x}$ to all points $\boldsymbol{x}_{i}$, $i=1,\ldots,n$. Set $a=\min\{f(\boldsymbol{x}_{i}),i=1,\ldots,n\}$, $b=\max\{f(\boldsymbol{x}_{i}),i=1,\ldots,n\}$. Then divide $[a,b]$ into $K$ equal length intervals with $p$ percentage overlapping ratio between any two adjacent intervals, denoted as $I_{j}=[a_{j},b_{j}]$, $j=1,\ldots,K$, as shown in the lower panel of Figure 1 (b). These intervals such that $[a,b]=\bigcup_{j=1}^{K}I_{j}$, $|b_{j}-a_{j+1}|=p|b_{1}-a_{1}|$, $j=1,\ldots,n-1$, where $a=a_{1}<b_{1}<a_{2}<\ldots<a_{n}<b_{n-1}<b_{n}=b$.
@@
Each interval $I_{j}$ is then pulled back to the original space through the inverse mapping $f^{-1}(I_{j}),j=1,...,K$. A clustering algorithm is applied to original data points fall into each $f^{-1}(I_{j})$, partitioning the points into $r_{j}$ disjoint clusters, represented as $C_{j,i}$, $i=1,\ldots,r_{j}$, such that $\bigcup_{r_{j}}C_{j,r_{j}}\subset f^{-1}(I_{j})$ ,where $j=1,...,K$. By pulling back all intervals, we can define a pull back cover $C=\{C_{1,1},...,C_{1,r_{1}},...,C_{K,1},...,C_{K,r_{K}}\}$. The Mapper graph is constructed from this cover, where each cluster in $C$ corresponds to a node in the graph. An edge is added between any two nodes if their corresponding clusters intersect, i.e., if $C_{i,j}\cap C_{i^{\prime},j^{\prime}}\neq\varnothing$. As shown in Figure 1 (c), the resulting Mapper graph accurately captures the shape of the data.
@@ 2.2 Soft Mapper
Traditional Mapper graphs are constructed through intervals based on the filtered data. The soft Mapper constructs Mapper graphs with a hidden assignment matrix without requiring fixed intervals (Oulhaj et al., 2024). The hidden assignment matrix is a ${n\times K}$ matrix, depicts the allocation relationship between given $n$ data points and $K$ groups. This matrix is denoted as $H=(H_{ij})_{n\times K}$, in which $H_{ij}=1$ if the $i$-th point belongs to $j$-th interval, otherwise $H_{ij}=0$, for $i=1,\ldots,n,~{}j=1,\ldots,K$. With a hidden assignment matrix $H$, a Mapper graph is constructed directly through the standard process of pulling back and clustering, which we refer to as a Mapper function, defined as follows.
@@
A Mapper function $\phi:H\mapsto G$, is a map from a hidden assignment matrix $H$ to a Mapper graph $G$. This function includes pulling back and clustering operations.
@@ 3.1 GMM Soft Mapper
By incorporating a Gaussian Mixture Model, we can naturally define a hidden assignment matrix $H$ through soft clustering, and a probability matrix $Q$ can be easily derived from the GMM parameters. We fit a GMM to the projected data $\{y_{1},...,y_{n}\}$. As a result, the $Q_{ij}$ represents the probability that data point $y_{i}$ is assigned to $j$-th class within the GMM framework.
@@
where $i=1,\dots,n,j=1,\dots,K$. This GMM-based assignment scheme imposes an implicit constraint on each row of the probability matrix $Q$, such that the sum of the probabilities for each data point equals one, $\sum_{j=1}^{K}Q_{ij}=1$. However, the independent Bernoulli distribution assignment scheme of $H_{ij}$ in the Soft Mapper may lead to extreme assignments, some points being unassigned to any clusters or a point is being assigned to all clusters. For instance, if $K=4$, data point $y_{i}$ might have an assignment probability vector of $Q_{i.}=[0.3,0.2,0.4,0.1]$, $i=1,\ldots,n$. Then the independent Bernoulli distribution assignment scheme may result an assignment vector $H_{i.}=[0,0,0,0]$, which means the data point $y_{i}$ is being unassigned to any clusters, or $H_{i.}=[1,1,1,1]$, which means the data point $y_{i}$ belongs to all clusters. The former case violates the Mapper algorithm requirement that each data point belongs to at least one cluster, and the extreme case of being unassigned to any clusters should be avoided. The latter case conflicts with that fact the a data point in the Mapper algorithms typically falls into, at most, two adjacent intervals. Therefore, we propose a GMM-multinomial soft Mapper approach. Instead of sampling each element of $H$ independent through the Bernoulli distributions, we sample each row of the assignment matrix $H$ through a multinomial distributionconcerning each row of the probability matrix $Q$. The multinomial distribution grantees that each point is assigned to at least to one cluster, and at most to $m$ clusters as follows.
@@
$H_{.}\sim{Multi}(m,~{}Q_{i.}),~{}i=1,\ldots,n.$
Here $H_{ij}$ may take values from $\{0,1,2\}$, and $H_{ij}\geq 1$ indicates that the $i$-th point is assigned to $j$-th interval, otherwise $H_{ij}=0$. For ease of notations, we denote $H\sim{Multi}(m,~{}Q)$ to indicate that each row of $H$ independently follows a multinomial distribution with event number parameter $m=2$ and event probabilities being each row of $Q$.
Our proposed soft Mapper samples each row $H_{i\cdot}$ independently from a multinomial distribution ${Multi}(2,Q_{i\cdot}),~{}i=1,\ldots,n.$. The total events $m$ of the multinomial distribution is set to $2$, guaranteeing that each data point can be assigned to at least one class and no more than two classes. This setting is reasonable, as in the Mapper algorithm, a data point typically falls within at least one interval and at most two adjacent intervals.
Our approach differs from the soft Mapper described in Oulhaj et al. (2024) in two key ways. Firstly, by using a GMM model, our $Q_{ij}$ can be specified as a function of the parameters $\boldsymbol{\theta}$. Thus, updating the GMM parameters $\boldsymbol{\theta}$ is equivalent to update $Q$. Whereas $Q$ in Oulhaj et al. (2024) was determined by the parameters in the filter functions. Secondly, we use an independent multinomial distribution assignment scheme to sample each row of the hidden assignment matrix $H$. The multinomial distribution avoids extreme cases where a data point is unassigned to any clusters or assigned to too many clusters.

### v2

@@ Abstract
The Mapper algorithm is an essential tool for visualizing complex, high dimensional data in topological data analysis (TDA) and has been widely used in biomedical research. It outputs a combinatorial graph whose structure encodes the shape of the data. However, the need for manual parameter tuning and fixed (implicit) intervals, along with fixed overlapping ratios may impede the performance of the standard Mapper algorithm. Variants of the standard Mapper algorithms have been developed to address these limitations, yet most of them still require manual tuning of parameters. Additionally, many of these variants, including the standard version found in the literature, were built within a deterministic framework and overlooked the uncertainty inherent in the data. To relax these limitations, in this work, we introduce a novel framework that implicitly represents intervals through a hidden assignment matrix, enabling automatic parameter optimization via stochastic gradient descent. In this work, we develop a soft Mapper framework based on a Gaussian mixture model(GMM) for flexible and implicit interval construction. We further illustrate the robustness of the soft Mapper algorithm by introducing the Mapper graph mode as a point estimation for the output graph. Moreover, a stochastic gradient descent algorithm with a specific topological loss function is proposed for optimizing parameters in the model. Both simulation and application studies demonstrate its effectiveness in capturing the underlying topological structures. In addition, the application to an RNA expression dataset obtained from the Mount Sinai/JJ Peters VA Medical Center Brain Bank (MSBB) successfully identifies a distinct subgroup of Alzheimer’s Disease. The implementation of our method is available at https://github.com/FarmerTao/Implicit-interval-Mapper.git.
@@ 1 Introduction
The Mapper algorithm is a powerful tool in topological data analysis (TDA) used to explore the shape of datasets. Initially introduced for 3D object recognition (Singh et al., 2007), the Mapper algorithm has demonstrated remarkable efficiency in extracting topological information across various data types. Its application spans various areas, notably in biomedical research (Skaf and Laubenbacher, 2022). For example, it successfully identified new breast cancer subgroups with superior survival rates (Nicolau et al., 2011). In single cell RNA-Seq analysis, a Mapper-based algorithm was used to study unbiased temporal transcriptional regulation (Rizvi et al., 2017). The variants of Mapper algorithms were also used to uncover higher-order structures of complex phenomics data (Kamruzzaman et al., 2021). Moreover, many efforts have been made to integrate the Mapper algorithm with other machine learning techniques. For instance, Bodnar et al. (2021) enhanced the performance of a graph neural network by incorporating the Mapper algorithm as a pooling operation. When combined with autoencoders, the Mapper algorithm can be used as a robust classifier (Cyranka et al., 2019), remedying the shortcomings of traditional convolutional neural networks, such as susceptibility to gradient-based attacks. Recently, a visualization method called ShapeVis which was inspired by the Mapper algorithm was proposed to produce more concise topological structure than standard Mapper graph (Kumari et al., 2020). Besides advances of the Mapper algorithms in applications, there has been some new progress in theory. Carrière et al. (2018) analyzed the Mapper algorithm from statistical perspective and introduced a novel method for parameter selection. Other studies have explored the algorithm’s convergence to Reeb graphs (Brown et al., 2021), and examined its topological properties (Dey et al., 2017).
@@
The standard Mapper algorithm generates a graph that captures the shape of the data by dividing the support of the filtered data into several overlapped intervals with a fixed length (Singh et al., 2007). Two major limitations of the standard Mapper algorithm are the requirements of fixed length of intervals and overlapping rate. Many algorithms were proposed to address these limitations. For instance, F-Mapper (Bui et al., 2020) utilized the fuzzy clustering algorithm to allow flexible interval partitioning by introducing an additional threshold parameter to control the degree of overlap. However, the introduction of the parameter also increased the algorithm’s computational complexity. In a different vein, the ensemble version of Mapper was proposed by Kang and Lim (2021); Fitzpatrick et al. (2023), which can recover the topology of datasets without parameter tuning. However, this method has a significant high computational cost due to the execution of multiple instances of the Mapper algorithm. Differently, the Ball Mapper (Dłotko, 2019) directly constructed cover on a given dataset, avoiding the difficulty of choosing filter functions. Nevertheless, its construction process is somewhat arbitrary and could inadvertently introduce extra parameters, increasing computational complexity. The D-Mapper, as introduced by Tao and Ge (2024), developed a probabilistic framework to construct intervals of the projected data based on a mixture distribution. This approach offers flexibility in interval construction. Additionally, the Soft Mapper, proposed by Oulhaj et al. (2024), a smoother version of the standard Mapper algorithm. This algorithm focused on the optimization of filter functions by minimizing the topological loss.
@@ 2.1 Mapper algorithm
The Mapper algorithm has been widely used to output a graph of a given dataset to represent its topological structure (Dey et al., 2017), the output graph can be regarded as a discrete version of the Reeb graph. The core concept of the Mapper algorithm is as follows. First, it projects a given high-dimensional dataset into a lower-dimensional space. Then, it constructs a reasonable cover for the projected data, pulls back the original data points according to each interval, and clusters the points within each preimage (or pullback set). Finally, each cluster is treated as a node of the output graph, and an edge is added between two clusters if they share any data points.
@@
Consider a dataset with $n$ data points, $\{\boldsymbol{x_{1}},\ldots,\boldsymbol{x_{n}}\}$, $\boldsymbol{x_{i}}\in\mathbb{R}^{d}$, $i=1,\ldots,n$. Let $\mathbb{X}_{n}$ be the support of this dataset. Denote $f(\cdot):\mathbb{X}^{n}\rightarrow\mathbb{R}^{m}$, $m<n$, a single-valued filter function, which maps each data point from $\mathbb{X}^{n}$ to a lower-dimensional space, $\mathbb{R}^{m}$ ($m\geq 1$). As in most of literature Bui et al. (2020); Kang and Lim (2021); Fitzpatrick et al. (2023), here $m$ is set to $1$. We use a simple example to illustrate the Mapper algorithm. The example dataset has a cross shape, as shown in Figure 1 (a). The Mapper algorithm first projects data onto a real line using a filter function, $f(\boldsymbol{x})=y,~{}\boldsymbol{x}\in\mathbb{R}^{2},~{}y\in\mathbb{R}$, as shown in the upper panel of Figure 1 (b), the filter function is set to $f(\boldsymbol{x})=\frac{1}{n}\sum_{i=1}^{n}||\boldsymbol{x}-\boldsymbol{x}_{i}||$, i.e. the mean distance of $\boldsymbol{x}$ to all points $\boldsymbol{x}_{i}$, $i=1,\ldots,n$. Set $a=\min\{f(\boldsymbol{x}_{i}),i=1,\ldots,n\}$, $b=\max\{f(\boldsymbol{x}_{i}),i=1,\ldots,n\}$. Then divide $[a,b]$ into $K$ equal length intervals with $p$ percentage overlapping ratio between any two adjacent intervals, denoted as $I_{j}=[a_{j},b_{j}]$, $j=1,\ldots,K$, as shown in the lower panel of Figure 1 (b). These intervals such that $[a,b]=\bigcup_{j=1}^{K}I_{j}$, $|b_{j}-a_{j+1}|=p|b_{1}-a_{1}|$, $j=1,\ldots,n-1$, where $a=a_{1}<b_{1}<a_{2}<\ldots<a_{n}<b_{n-1}<b_{n}=b$.
@@
Each interval $I_{j}$ is then pulled back to the original space through the inverse mapping $f^{-1}(I_{j}),j=1,...,K$. A clustering algorithm is applied to original data points fall into each $f^{-1}(I_{j})$, partitioning the points into $r_{j}$ disjoint clusters, represented as $C_{j,i}$, $i=1,\ldots,r_{j}$, such that $\bigcup_{r_{j}}C_{j,r_{j}}\subset f^{-1}(I_{j})$ ,where $j=1,...,K$. We then obtain the pullback cover, $C=\{C_{1,1},...,C_{1,r_{1}},...,C_{K,1},...,C_{K,r_{K}}\}$. The Mapper graph is constructed from this cover, where each cluster in $C$ corresponds to a node in the graph. An edge is added between any two nodes if their corresponding clusters intersect, i.e., if $C_{i,j}\cap C_{i^{\prime},j^{\prime}}\neq\varnothing$. As shown in Figure 1 (c), the resulting Mapper graph accurately captures the shape of the data.
@@ 2.2 Soft Mapper
Traditional Mapper graphs are constructed through intervals based on the filtered data. The soft Mapper constructs Mapper graphs with a hidden assignment matrix without requiring fixed intervals (Oulhaj et al., 2024). The hidden assignment matrix, a ${n\times K}$ matrix, depicts the allocation relationship between given $n$ data points and $K$ groups (i.e. the implicit $K$ intervals). This matrix is denoted as $H=(H_{ij})_{n\times K}$, in which $H_{ij}=1$ if the $i$-th point belongs to $j$-th group (implicit interval), otherwise $H_{ij}=0$, for $i=1,\ldots,n,~{}j=1,\ldots,K$. With a hidden assignment matrix $H$, a Mapper graph is constructed directly through the standard process of pulling back and clustering, which we refer to as a Mapper function, defined as follows.
@@
A Mapper function $\phi:H\mapsto G$, is a map from a hidden assignment matrix $H$ to a Mapper graph $G$. The function is defined by pulling back and clustering operations in the standard Mapper algorithm.
@@ 3.1 GMM Soft Mapper
In the standard Mapper algorithm, a cover is constructed on the projected data, and a cover consists of several intervals and two adjacent intervals overlap. These intervals can be constructed by partitioning the projected data into multiple groups. Then, pull back the data points by allocating the corresponding projected value into each interval. Analogous, this allocating process of projected data points into intervals of a cover is similar to the allocating process of labels for a mixture probability model. In this work, we focus on a Gaussian mixture model (GMM) to fit the projected data due to its simplicity and flexibility. By incorporating a GMM, we can naturally define a hidden assignment matrix $H$ through soft clustering, and a probability matrix $Q$ can be easily derived from the GMM parameters. We fit a GMM to the projected data $\{y_{1},...,y_{n}\}$. As a result, the $Q_{ij}$ represents the probability that data point $y_{i}$ is assigned to $j$-th class within the GMM framework.
@@
where $i=1,\dots,n,j=1,\dots,K$.
This GMM-based assignment scheme imposes an implicit constraint on each row of the probability matrix $Q$, such that the sum of the probabilities for each data point equals one, $\sum_{j=1}^{K}Q_{ij}=1$. However, the independent Bernoulli distribution assignment scheme of $H_{ij}$ in the Soft Mapper may lead to extreme assignments, some points being unassigned to any clusters or a point is being assigned to all clusters. For instance, if $K=4$, data point $y_{i}$ might have an assignment probability vector of $Q_{i.}=[0.3,0.2,0.4,0.1]$, $i=1,\ldots,n$. Then the independent Bernoulli distribution assignment scheme may result an assignment vector $H_{i.}=[0,0,0,0]$, which means the data point $y_{i}$ is being unassigned to any clusters, or $H_{i.}=[1,1,1,1]$, which means the data point $y_{i}$ belongs to all clusters. The former case violates the Mapper algorithm requirement that each data point belongs to at least one cluster, and the extreme case of being unassigned to any clusters should be avoided. The latter case conflicts with that fact the a data point in the Mapper algorithms typically falls into, at most, two adjacent intervals. Therefore, we propose a GMM-multinomial soft Mapper approach. Instead of sampling each element of $H$ independently through the Bernoulli distributions, we sample each row of the assignment matrix $H$ through a multinomial distributionconcerning each row of the probability matrix $Q$. The multinomial distribution grantees that each point is assigned to at least to one cluster, and at most to $m$ clusters as follows.
@@
$H_{i\cdot}\sim{Multi}(m,~{}Q_{i.}),~{}i=1,\ldots,n.$
Here $H_{ij}$ may take values from $\{0,1,2\}$, and $H_{ij}\geq 1$ indicates that the $i$-th point is assigned to $j$-th implicit interval (group), otherwise $H_{ij}=0$. For ease of notations, we denote $H\sim{Multi}(m,~{}Q)$ to indicate that each row of $H$ independently follows a multinomial distribution with event number parameter $m=2$ and event probabilities being each row of $Q$.
Our proposed soft Mapper samples each row $H_{i\cdot}$ independently from a multinomial distribution ${Multi}(2,Q_{i\cdot}),~{}i=1,\ldots,n$. The total events $m$ of the multinomial distribution is set to $2$, guaranteeing that each data point must be assigned to at least one group and no more than two groups. This setting is reasonable, as in the Mapper algorithm, a data point typically falls within at least one group and at most two groups.
Our approach differs from the soft Mapper described in Oulhaj et al. (2024) in two key ways. Firstly, by using the GMM, our $Q_{ij}$ can be specified as a function of the parameters $\boldsymbol{\theta}$. Thus, updating the GMM parameters $\boldsymbol{\theta}$ is equivalent to updating $Q$. In contrast, $Q$ in Oulhaj et al. (2024) was determined by the parameters in the filter functions. Secondly, we use an independent multinomial distribution assignment scheme to sample each row of the hidden assignment matrix $H$. The multinomial distribution avoids extreme cases where a data point is unassigned to any clusters or assigned to too many clusters.